OpenStudy (republican31):

y + 5 = 1/2(x + 2) in Standard Form for Graphing. I think it would go like this: -2y - 10 = 1(x + 2) -2y - 10 = x + 2 ? The actual answer is x – 2y = 8 but I don't understand how?

OpenStudy (kearsten):

solve for x?

OpenStudy (republican31):

No this is converting it to standard form for graphing. Sorry forgot to mention that.

OpenStudy (republican31):

Please look at edits by the way.

OpenStudy (kearsten):

im not very good at math but if you type it in search all kinds of solutions and help pages pop up hope they help bcu i can give you the answer but the steps are what im not very good at sry

OpenStudy (republican31):

Alright. The problem is I have the answer from the lesson (it's a practice question) but I don't understand how that is the answer: x – 2y = 8. Thank you for your help!

Vocaloid (vocaloid):

start with: y + 5 = 1/2(x + 2) multiply both sides by 2 gives us: 2y + 10 = (x+2) subtract 2 from both sides gives us: 2y + 8 = x subtract 2y from both sides gives us 8 = x - 2y flip the equation around gives us: x - 2y = 8

OpenStudy (republican31):

THANK YOU! I was originally trying to move the 8 over to the 2. And "originally" means before I tried dividing by -2, thinking I might have been supposed to for moving it to the other side... I don't know. Anyways, thank you!